Calculate the heat energy released when 25.1g of liquid mercury at 25.00 °C is converted to solid mercury at its melting point.
1
Expert's answer
2020-02-19T07:58:37-0500


We will calculate the following steps:


Liquid mercury releases part of its internal energy into the environment when it cools and then hardens.


1.Mercury cooling:

Q = mcΔT= 25.1 g*0.140 J/g·°C*(-38.83° C - 25.00° C) = -224J

2.Crystallization of mercury:

Q = mL = 25.1 g*(-11.6 J/g) = -291.16 J

Q=-224 J-291.16=515.16 J